OpenStudy (anonymous):

A hotel lobby can seat 19 people at chairs and loveseats. The lobby has a total of 12 pieces of furniture for seating. Assuming a chair seats 1 person and a loveseat seats 2, how many loveseats are in the lobby? 4 5 6 7

OpenStudy (anonymous):

we can use algebra, but it is much easier to guess and check for this one lets guess 6 since it is usually C if there are 6 love seats then there are 6 chairs that would seat 12 in the love seats and 6 in the chairs for a total of 18, but you want 19 so C is not the answer

OpenStudy (anonymous):

try D

OpenStudy (anonymous):

i got 14 for d

hero (hero):

x = chairs y = loveseats x + y = 12 x + 2y = 19 Just use the algebra

OpenStudy (anonymous):

i got 6

hero (hero):

Nope, not correct

OpenStudy (anonymous):

cause 12 dived 2

hero (hero):

Let me post the steps

OpenStudy (anonymous):

we checked that 6 did not work

hero (hero):

x + y = 12 x + 2y = 19 Mulitiply the first equation by -1 to get -x - y = -12 x + 2y = 19 Add the two equations together to get: y = 7

OpenStudy (anonymous):

6 love seats seats \(6\times 2=12\) people if there are 6 love seats there are also 6 chairs (since the total is 12) and they each seat only one person and 12 + 6 = 18 but you want 19

OpenStudy (anonymous):

so 7

hero (hero):

Well y was defined as loveseats so what do you think?

OpenStudy (anonymous):

i think its 7 cause 7+12=19

OpenStudy (anonymous):

i think